OpenStudy (anonymous):

what is the rationalizing factor of root 3+1????????? pls ans fast

OpenStudy (anonymous):

\[\sqrt3+1\] i guess it would be the \(\sqrt3-1\) although i have never heard it called that before

OpenStudy (anonymous):

"rationalizing factor"?? learn something new every day

OpenStudy (anonymous):

Do you mean something along these lines? \[\frac{x}{(\sqrt{3}+1)} \] And you need to rationalize your denominator? What @satellite73 sent you is the correct answer if that is the case

OpenStudy (anonymous):

It is called that because it is the factor used to rationalize the denominator when involving complex factors. And radical factors.

OpenStudy (anonymous):

Meaning something like \[\frac{x}{\sqrt{3}+1}=\frac{x(\sqrt{3}-1)}{(\sqrt{3}+1)(\sqrt{3}-1)}=\frac{x(\sqrt{3}-1)}{2}\]?

OpenStudy (anonymous):

Yep That is exactly what it is. Or at least what I am talking about.

OpenStudy (anonymous):

Ok :) Another term learned, thanks.

OpenStudy (anonymous):

thnx ppl